# Exploiting Agent Symmetries for Performance Analysis of Distributed Optimization Methods

### v1

@@ 2.2 Representation of consensus steps in PEP
[…]
$\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}=\left\{\!\begin{aligned} &W^{T}=W\\ W\in\mathbb{R}^{n\times n}:\quad&\lambda_{1}(W)=1,\leavevmode\nobreak\ v_{1}(W)=\mathbf{1}/{\scriptstyle\sqrt{n}}\\ &\lambda^{-}\leq\lambda_{n}(W)\leq\dots\leq\lambda_{2}(W)\leq\lambda^{+}\end{aligned}\right\}.$ (20)
Such a set is often used to derive theoretical bounds on distributed algorithms [20]. You may notice that it is not restricted to non-negative matrices, even if the literature often assumes non-negativity through doubly-stochasticity. We motivate this choice by two arguments. Firstly, non-negativity is not needed for convergence of pure consensus steps [33]. Secondly, among results using doubly-stochasticity, we found no results exploiting the non-negativity of the matrix, see e.g. [20, 24, 15]. Such results are thus about generalized doubly-stochastic matrices [9], which refers to matrices whose rows and columns sum to one, as the ones contained in $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$.
This section first shows why there is no Gram representation of the consensus steps using a matrix from the set $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, which would have allowed integrating them in an SDP PEP (see Proposition 2). Secondly, we present a relaxation to obtain a Gram-representable description of such consensus steps and we characterize the set of matrices that are tightly described using this description.
@@ 2.2.1 The intractable quest for tightness
It would be ideal to find necessary and sufficient conditions for a set of pairs $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}$ to be $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable, see Definition 4. Moreover, the conditions should be Gram-representable, i.e. linear in the scalar products of $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$, to be able to use them in an SDP PEP formulation (see Proposition 2).
Let $I$ be a set of indices of consensus steps. A set of pairs $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ is $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable if,
$\exists W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathbf{y}^{k}=(W\otimes I_{d})\mathbf{x}^{k}\quad\text{ for all $k\in I.$}$ (21)
In this subsection, we will show, by a non-convexity argument, that there are no tight conditions for $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolability that are Gram-representable.
@@
We define $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ as the set of symmetric and positive semi-definite Gram matrices of the form
@@
such that the set of pairs $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k=1,\dots,t}$ is $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable.
In practice, in the PEPs, Gram matrices can involve other variables than those of the consensus (e.g. gradients). We can therefore see $G_{c}$ as a submatrix of the Gram matrix with all the scalar products related to consensus steps.
@@
[…]
@@
We build two Gram matrices $G_{1},G_{2}\in\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ and show that their combination $G_{3}=\frac{1}{2}(G_{1}+G_{2})$ is not in $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$. We build a counter-example for one consensus step, i.e. for $t=1$, and then we explain why it can also be used for $t>1$. In the set $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, there is no constraint on the rank $d$ of the matrices, so we choose to build $G_{1}$ and $G_{2}$ with rank $d=1$, as follows
let $\text{let }\quad\mathbf{x}=\small{\begin{bmatrix}n\\ 0\\ \vdots\\ 0\end{bmatrix}}\in\mathbb{R}^{n},$ $[W_{1}]_{ij}$ =$=\begin{cases}\frac{1+(n-1)\lambda^{-}}{n}&\text{for $i=j$}\\ \frac{1-\lambda^{-}}{n}&\text{for $i\neq j$}\end{cases}\quad$ $[W_{2}]_{ij}$ =$=\begin{cases}\frac{1+(n-1)\lambda^{+}}{n}&\text{for $i=j$}\\ \frac{1-\lambda^{+}}{n}&\text{for $i\neq j$}\end{cases},$ (31)
and $\mathbf{y}_{1}$ $=W_{1}\mathbf{x}=\footnotesize{\begin{bmatrix}1+(n-1)\lambda^{-}\\ 1-\lambda^{-}\\[-2.84526pt] \vdots\\ 1-\lambda^{-}\end{bmatrix}},$ $\mathbf{y}_{2}$ $=W_{2}\mathbf{x}=\footnotesize{\begin{bmatrix}1+(n-1)\lambda^{+}\\ 1-\lambda^{+}\\[-2.84526pt] \vdots\\ 1-\lambda^{+}\end{bmatrix}}$ (40)
we build $G_{1}$ and $G_{2}$ as $G_{1}$ $=\begin{bmatrix}\mathbf{x}\mathbf{x}^{T}&\mathbf{x}\mathbf{y}_{1}^{T}\\ \mathbf{y}_{1}\mathbf{x}^{T}&\mathbf{y}_{1}\mathbf{y}_{1}^{T}\end{bmatrix},\quad$ $G_{2}$ $=\begin{bmatrix}\mathbf{x}\mathbf{x}^{T}&\mathbf{x}\mathbf{y}_{2}^{T}\\ \mathbf{y}_{2}\mathbf{x}^{T}&\mathbf{y}_{2}\mathbf{y}_{2}^{T}\end{bmatrix}.$ (45)
[…]
$G_{3}=\frac{1}{2}(G_{1}+G_{2})=\begin{bmatrix}\mathbf{x}\mathbf{x}^{T}&\frac{1}{2}\mathbf{x}(\mathbf{y}_{1}+\mathbf{y}_{2})^{T}\\ \frac{1}{2}(\mathbf{y}_{1}+\mathbf{y}_{2})\mathbf{x}^{T}&\frac{1}{2}(\mathbf{y}_{1}\mathbf{y}_{1}^{T}+\mathbf{y}_{2}\mathbf{y}_{2}^{T})\end{bmatrix}=\begin{bmatrix}P_{X_{3}}^{T}P_{X_{3}}&P_{X_{3}}^{T}P_{Y_{3}}\\ P_{Y_{3}}^{T}P_{X_{3}}&P_{Y_{3}}^{T}P_{Y_{3}}\end{bmatrix}.$ (46)
@@
[…]
When we consider multiple consensus steps using the same matrix, i.e. when $t>1$, the counter-example developed above for $t=1$ can still be chosen for one of the steps, and the Gram matrices $G_{1}$, $G_{2}$, and $G_{3}$ would then be submatrices of the Gram matrices in $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, showing the non-convexity of $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ for $t>1$ as well. ∎
@@ 2.2.2 A relaxed formulation
In previous work [5], necessary constraints have been used to represent the effect of consensus steps in PEPs. These constraints describe a convex set of Gram matrices that contains the non-convex set $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, and can therefore be exploited in PEP to derive upper bounds on worst-case performance of distributed algorithms. However, the precise meaning of these constraints was not known. In this subsection, we show that these constraints correspond to tight interpolation constraints for a larger class of steps, containing the consensus step as a particular case.
In what follows, we assume $X,Y\in\mathbb{R}^{nd\times t}$ to be the $t$ points linked by a symmetric linear operator $Y=MX$ that preserves the projection of their columns onto a given subspace $\mathcal{B}$, denoted respectively $X_{\mathbin{\!/\mkern-5.0mu/\!}}$ and $Y_{\mathbin{\!/\mkern-5.0mu/\!}}$, i.e. $Y_{\mathbin{\!/\mkern-5.0mu/\!}}=X_{\mathbin{\!/\mkern-5.0mu/\!}}$. We consider the decomposition of matrices of points $X$ and $Y$ in two orthogonal terms:
$X=X_{\mathbin{\!/\mkern-5.0mu/\!}}+X_{\perp}\text{ and }Y=Y_{\mathbin{\!/\mkern-5.0mu/\!}}+Y_{\perp},$
where $Y_{\perp},X_{\perp}\in\mathbb{R}^{nd\times t}$ are obtained by projecting each column of $X$ and $Y$ on the orthogonal subspace $\mathcal{B}^{\perp}$. Therefore, the described matrix step acts on both terms as follows:
$Y=MX=M(X_{\mathbin{\!/\mkern-5.0mu/\!}}+X_{\perp})=X_{\mathbin{\!/\mkern-5.0mu/\!}}+MX_{\perp}\qquad\Leftrightarrow\qquad Y_{\mathbin{\!/\mkern-5.0mu/\!}}=X_{\mathbin{\!/\mkern-5.0mu/\!}}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ Y_{\perp}=MX_{\perp}.$
Consensus steps (6) with symmetric and generalized doubly-stochastic averaging matrices (20) are particular cases of these general matrix steps where the preserved subspace is the consensus subspace (see Definition 6). Indeed, a consensus step always preserves the agent average; which corresponds to the projection on the consensus subspace: $\overline{Y}=\overline{X}$. Moreover, consensus steps (6) require a specific shape for matrix $M=W\otimes I_{d}$, to ensure it applies the same block matrix $W\in\mathbb{R}^{n\times n}$ independently to each dimension of the variables.
@@
$\mathcal{C}=\left\{\mathbf{x}\in\mathbb{R}^{nd}\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{1}=\dots=x_{n}\in\mathbb{R}^{d}\right\}.\vspace{-1mm}$
@@
$\mathcal{C}^{\perp}=\left\{\mathbf{x}\in\mathbb{R}^{nd}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}^{T}\mathbf{y}=0,\text{ for all }\mathbf{y}\in\mathcal{C}\right\}.$
Formally, we assume the matrices $M$ to be part of the following set
$\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{B}}=\left\{\!\begin{aligned} &M^{T}=M\\ \leavevmode\nobreak\ M\in\mathbb{R}^{nd\times nd}:\quad&\lambda_{1}(M)=\dots=\lambda_{|\mathcal{B}|}(M)=1,\leavevmode\nobreak\ \text{with eigenvectors forming a basis of $\mathcal{B}$}\\ &\lambda^{-}\leq\lambda_{nd}(M)\leq\dots\leq\lambda_{|\mathcal{B}|+1}(M)\leq\lambda^{+}\end{aligned}\right\}.$ (48)
As in $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, we impose symmetry of the matrix, which allows working with real eigenvalues. Extension to non-symmetric averaging matrices would require defining matrix sets $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ and $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ based on singular values. We now provide tight interpolation constraints for this set $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ of symmetric matrices.
Definition 7 ($\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{B}}$-interpolability).
Let $I$ be a finite set of indices of matrix steps and $S$ be a subspace of $\mathbb{R}^{nd}$. A set of pairs of points $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ is $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{B}}$-interpolable if,
@@
$\exists M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{B}}\leavevmode\nobreak\ :\leavevmode\nobreak\ \mathbf{y}^{k}=M\mathbf{x}^{k}\quad\text{ for all $k\in I$.}$ (49)
In this definition, the pairs of points $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ are all linked by the same matrix $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$. Thus, settings or algorithms using different matrices work with different sets of pairs of points. This definition is related to the interpolation of a symmetric linear operator, developed in [1]. In this paper, we focus on linear operators preserving a given subset $\mathcal{B}$, which can be used to represent consensus steps. Based on results from [5] and [1], we give necessary and sufficient conditions to have $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{B}}$-interpolability of a set of pairs of points.
Theorem 4 ($\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{B}}$ interpolation constraints).
@@
Let $I$ be a finite set of indices of matrix steps and $S$ a subspace of $\mathbb{R}^{nd}$. A set of pairs of points $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ is $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}^{\mathcal{B}}$-interpolable if, and only if,

### v2

@@ 2.2 Representation of consensus steps in PEP
In PEP for a decentralized optimization algorithm $\mathcal{A}\in\mathcal{A}_{D}$, with $n\geq 2$ agents, we would like to find Gram-representable constraints to represent all the pairs of points $\{\mathbf{x}^{k},\mathbf{y}^{k}\}$ such that
$\mathbf{y}^{k}=(W\otimes I_{d})\mathbf{x}^{k},\quad\text{for $k=1,\dots,t$},\qquad\text{with }\penalty 10000\ W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}},$ (21)
where $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ is the following set of symmetric averaging matrices, with fixed bounds $\lambda^{-},\lambda^{+}\in(-1,1)$ on the eigenvalues (except for $\lambda_{1}=1$) :
$\hskip-8.53581pt\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}=\left\{\!\begin{aligned} &W^{T}=W\\ W\in\mathbb{R}^{n\times n}:\quad&\lambda_{1}(W)=1,\penalty 10000\ v_{1}(W)=\mathbf{1}/{\scriptstyle\sqrt{n}}\\ &\lambda^{-}\leq\lambda_{n}(W)\leq\dots\leq\lambda_{2}(W)\leq\lambda^{+}\end{aligned}\right\}.$ (22)
Such a set is often used to derive theoretical bounds on distributed algorithms [21]. You may notice that it is not restricted to non-negative matrices, even if the literature often assumes non-negativity through doubly-stochasticity. We motivate this choice by two arguments. Firstly, non-negativity is not needed for the convergence of pure consensus steps [34]. Secondly, among results using doubly-stochasticity, we found no results exploiting the non-negativity of the matrix, see e.g. [21, 25, 16]. Such results are thus about generalized doubly-stochastic matrices [10], which refers to matrices whose rows and columns sum to one, as the ones contained in $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$.
A set of pairs of points $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}$ satisfying the consensus steps constraints (21) is called $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable.
Let $I$ be a finite set of indices. A set of pairs $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ is $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable if,
$\exists W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}\penalty 10000\ :\penalty 10000\ \mathbf{y}^{k}=(W\otimes I_{d})\mathbf{x}^{k}\quad\text{ for all $k\in I.$}$ (23)
In this definition, the pairs of points $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ are all linked by the same averaging matrix $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$. Thus, settings or algorithms using different averaging matrices work with different sets of pairs of points.
This section first shows why there is no Gram-representable constraints to characterize $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolability of a set of pairs of points $\{\mathbf{x}^{k},\mathbf{y}^{k}\}$ involved in consensus steps of the form of (21), preventing its tight representation in an SDP PEP formulation. Then we present a relaxed Gram-representable description of $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable pairs of points and characterize the points allowed in this relaxed description. Finally, we also explain how it can be used to build the spectral agent-dependent PEP formulation, which provides worst-case guarantees for a decentralized algorithm, that are valid for any averaging matrix from $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$.
@@ 2.2.1 The intractable quest for tightness
It would be ideal to find necessary and sufficient conditions for a set of pairs $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}$ to be $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable, see Definition 4. Moreover, the conditions should be Gram-representable, i.e., linear in the scalar products of $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$, to be able to use them in an SDP PEP formulation (see Proposition 2). In this subsection, we will show, by a non-convexity argument, that there are no tight conditions for $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolability that are Gram-representable.
@@
We define $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ as the set of symmetric and positive semidefinite Gram matrices of the form
@@
such that the set of pairs $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k=1,\dots,t}$ is $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable, in the sense of Definition 4.
In practice, in the PEPs, the Gram matrices can involve other variables than those of the consensus (e.g., gradients). We can therefore consider $G_{c}$ as a submatrix of the full Gram matrix, which contains only the scalar products related to consensus steps.
@@
Let $\lambda^{-}<\lambda^{+}\in(-1,1)$. The set $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ is non-convex.
@@
We build two Gram matrices $G_{1},G_{2}\in\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ and show that their combination $G_{3}=\frac{1}{2}(G_{1}+G_{2})$ is not in $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$. We build a counter-example for one consensus step, i.e., for $t=1$, and then we explain why it can also be used for $t>1$. In the set $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, there is no constraint on the rank $d$ of the matrices, so we choose to build $G_{1}$ and $G_{2}$ with rank $d=1$, as follows
let $\mathbf{x}$ $=\small{\begin{bmatrix}n&0&\dots&0\end{bmatrix}^{T}}\in\mathbb{R}^{n},$ (27)
$[W_{1}]_{ij}$ $=\begin{cases}\frac{1+(n-1)\lambda^{-}}{n}&\text{for $i=j$}\\ \frac{1-\lambda^{-}}{n}&\text{for $i\neq j$}\end{cases},$ $[W_{2}]_{ij}$ $=\begin{cases}\frac{1+(n-1)\lambda^{+}}{n}&\text{for $i=j$}\\ \frac{1-\lambda^{+}}{n}&\text{for $i\neq j$}\end{cases},$ (28)
and $\mathbf{y}_{1}$ $=W_{1}\mathbf{x}=\footnotesize{\begin{bmatrix}1+(n-1)\lambda^{-}\\ 1-\lambda^{-}\\[-2.84526pt] \vdots\\ 1-\lambda^{-}\end{bmatrix}}$, $\hskip-28.45274pt\mathbf{y}_{2}=W_{2}\mathbf{x}=\footnotesize{\begin{bmatrix}1+(n-1)\lambda^{+}\\ 1-\lambda^{+}\\[-2.84526pt] \vdots\\ 1-\lambda^{+}\end{bmatrix}}$, (29)
we build $G_{1}$ and $G_{2}$ as
$G_{1}$ $=\begin{bmatrix}\mathbf{x}\mathbf{x}^{T}&\mathbf{x}\mathbf{y}_{1}^{T}\\ \mathbf{y}_{1}\mathbf{x}^{T}&\mathbf{y}_{1}\mathbf{y}_{1}^{T}\end{bmatrix}$$\quad$ $G_{2}$ $=\begin{bmatrix}\mathbf{x}\mathbf{x}^{T}&\mathbf{x}\mathbf{y}_{2}^{T}\\ \mathbf{y}_{2}\mathbf{x}^{T}&\mathbf{y}_{2}\mathbf{y}_{2}^{T}\end{bmatrix}$. (30)
The eigenvalues of $W_{1}$ are $\{\lambda^{-},\dots,\lambda^{-},1\}$, and those of $W_{2}$ are $\{\lambda^{+},\dots,\lambda^{+},1\}$, so that $W_{1},W_{2}\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$. Therefore, by construction, both pairs $(\mathbf{x},\mathbf{y}_{1})$ and $(\mathbf{x},\mathbf{y}_{2})$ are $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable, and so $G_{1}$ and $G_{2}$ are in $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, and can be combined as $G_{3}=\frac{1}{2}(G_{1}+G_{2})$,
$G_{3}=\begin{bmatrix}\mathbf{x}\mathbf{x}^{T}&\frac{1}{2}\mathbf{x}(\mathbf{y}_{1}+\mathbf{y}_{2})^{T}\\ \frac{1}{2}(\mathbf{y}_{1}+\mathbf{y}_{2})\mathbf{x}^{T}&\frac{1}{2}(\mathbf{y}_{1}\mathbf{y}_{1}^{T}+\mathbf{y}_{2}\mathbf{y}_{2}^{T})\end{bmatrix}=\begin{bmatrix}P_{X_{3}}^{T}P_{X_{3}}&P_{X_{3}}^{T}P_{Y_{3}}\\ P_{Y_{3}}^{T}P_{X_{3}}&P_{Y_{3}}^{T}P_{Y_{3}}\end{bmatrix}$. (31)
@@
Then, the block $P_{X_{3}}^{T}P_{X_{3}}=\mathbf{x}\mathbf{x}^{T}$ has rank one, because $\mathbf{x}\in\mathbb{R}^{n}$, and thus the block $P_{Y_{3}}^{T}P_{Y_{3}}$, which can be expressed as $P_{Y_{3}}^{T}P_{Y_{3}}=W_{3}P_{X_{3}}^{T}P_{X_{3}}W_{3}$ using (32), also has rank one [^1] This is due to $\mathrm{rank}(AB)\leq\min(\mathrm{rank}(A),\mathrm{rank}(B))$.. However, the block is also defined as $P_{Y_{3}}^{T}P_{Y_{3}}=\frac{1}{2}(\mathbf{y}_{1}\mathbf{y}_{1}^{T}+\mathbf{y}_{2}\mathbf{y}_{2}^{T})$ from (31), which has rank 2 for any $\lambda^{-}\neq\lambda^{+}\in(-1,1)$, because vectors $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$ (29) are linearly independent in that case. By this contradiction argument, we have proved that $G_{3}=\frac{1}{2}(G_{1}+G_{2})$ is not in $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, meaning that this set is not convex for $t=1$.
When we consider multiple consensus steps using the same matrix, i.e., when $t>1$, the counter-example developed above for $t=1$ can still be chosen for one of the steps, and the Gram matrices $G_{1}$, $G_{2}$, and $G_{3}$ would then be submatrices of the Gram matrices in $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, showing the non-convexity of $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ for $t>1$ as well.
Proof of Theorem 3.
By definition, Gram-representable constraints define a convex subset of symmetric and semidefinite Gram-matrices. However, by Lemma 3.1, we know that set $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ of Gram-matrices such that the pairs of points $\{\mathbf{x}^{k},\mathbf{y}^{k}\}$ are $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable is a non-convex. Therefore, we know that there exists no Gram-representable constraints that tightly describe all the points $\{\mathbf{x}^{k},\mathbf{y}^{k}\}$ that are $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable. ∎
@@ 2.2.2 A relaxed formulation
In previous work [5], necessary constraints have been used to represent the effect of consensus steps in PEPs. These constraints describe a convex set of Gram matrices that contains the non-convex set $\mathcal{G}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, and can therefore be exploited in PEP to derive upper bounds on the worst-case performance of distributed algorithms. However, the precise meaning of these constraints was not known. In this subsection, we show that these constraints correspond to tight interpolation constraints for a larger class of steps, containing the consensus step as a particular case.
Generalized consensus steps
The consensus steps (21) can be seen as a linear operation on vectors $\mathbf{x}^{k}\in\mathbb{R}^{nd}$, with a matrix $M\in\mathbb{R}^{nd\times nd}$ that has a specific Kronecker structure :
$\mathbf{y}^{k}=M\mathbf{x}^{k},\quad\text{for $k=1,\dots,t$}\qquad\text{with }\penalty 10000\ M=W\otimes I_{d}\penalty 10000\ \text{ and }\penalty 10000\ W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}.$ (33)
By definition, matrix $M$ has the same eigenvalues as $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ (22), with multiplicity $d$ times larger. Therefore, it has $d$ eigenvalues equal to $1$, and their associated eigenvectors form a basis of the consensus subspace $\mathcal{C}$.
@@
$\mathcal{C}=\left\{\mathbf{x}\in\mathbb{R}^{nd}\penalty 10000\ |\penalty 10000\ x_{1}=\dots=x_{n}\in\mathbb{R}^{d}\right\}.$ (34)
@@
$\mathcal{C}^{\perp}=\left\{\mathbf{x}\in\mathbb{R}^{nd}\penalty 10000\ |\penalty 10000\ \mathbf{x}^{T}\mathbf{y}=0,\text{ for all }\mathbf{y}\in\mathcal{C}\right\}.$ (36)
We consider generalized consensus steps where the Kronecker structure of $M$ is relaxed, so that the vectors $\mathbf{x}^{k}$ are multiplied by a full matrix with similar properties :
@@
$\mathbf{y}^{k}=M\mathbf{x}^{k},\quad\text{for $k=1,\dots,t$},\qquad\text{with }\penalty 10000\ M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}},$ (38)
where $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ is a set of symmetric matrices that have the same properties as $W\otimes I_{d}$ with $W\in\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, but without constraint on their structure. In particular, matrices in $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ are still symmetric, generalized doubly stochastic and with the same range of eigenvalues. The set $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ is formally defined as follows, for a given range of eigenvalue $\lambda^{-},\lambda^{+}\in(-1,1)$, $\lambda^{-}\leq\lambda^{+}$ :
$\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}=\left\{\!\begin{aligned} &M^{T}=M\\ \penalty 10000\ M\in\mathbb{R}^{nd\times nd}:\quad&\lambda_{1}(M)=\dots=\lambda_{|\mathcal{C}|}(M)=1,\\[-5.69054pt] &\text{\small{with eigenvectors forming a basis of $\mathcal{C}$}}\\ &\lambda^{-}\leq\lambda_{nd}(M)\leq\dots\leq\lambda_{|\mathcal{C}|+1}(M)\leq\lambda^{+}\end{aligned}\right\},$ (40)
We say that a matrix $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ preserves the consensus subspace $\mathcal{C}$ as we have $M\mathbf{v}=\mathbf{v}$ for any $\mathbf{v}\in\mathcal{C}$. Relaxation of the Kronecker structure in a consensus step allows different weights to be applied to each component of the agent vectors, and different components to be mixed.
Interpolation constraints for $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$
We provide tight interpolation constraints on a set of pairs of points $\{\mathbf{x}^{k},\mathbf{y}^{k}\}$ for such generalized consensus steps (38), hence providing necessary constraints for describing classical consensus steps (21). Interpolation constraints are necessary and sufficient conditions for a set of pairs of points $\{\mathbf{x}^{k},\mathbf{y}^{k}\}$ to be interpolable by a matrix $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$.
Definition 7 ($\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolability).
@@
Let $I$ be a finite set of indices. A set of pairs of points $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ is $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable if,
$\exists M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}\penalty 10000\ :\penalty 10000\ \mathbf{y}^{k}=M\mathbf{x}^{k}\quad\text{ for all $k\in I$.}$ (41)
This definition is related to the interpolation of a symmetric linear operator, developed in [1]. In this work, we focus on linear operators preserving a given subspace, namely the consensus subspace $\mathcal{C}$. Based on results from [5] and [1], we give necessary and sufficient conditions to have $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolability of a set of pairs of points. Before stating these interpolation conditions, we first introduce a few notions. We aggregate all the vectors $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ in matrices $X,Y\in\mathbb{R}^{nd\times t}$ :
$X=[\mathbf{x}^{1}\dots\mathbf{x}^{t}]\quad\text{ and }\quad Y=[\mathbf{y}^{1}\dots\mathbf{y}^{t}]$ (43)
We also consider the decomposition of matrices $X$ and $Y$ in two orthogonal terms :
$X=X_{\mathbin{\!/\mkern-5.0mu/\!}}+X_{\perp}\text{ and }Y=Y_{\mathbin{\!/\mkern-5.0mu/\!}}+Y_{\perp},$ (45)
where $X_{\mathbin{\!/\mkern-5.0mu/\!}},Y_{\mathbin{\!/\mkern-5.0mu/\!}}\in\mathbb{R}^{nd\times t}$ and $Y_{\perp},X_{\perp}\in\mathbb{R}^{nd\times t}$ are obtained by projecting each column of $X$ and $Y$ respectively on the consensus subspace $\mathcal{C}$ and on its orthogonal subspace $\mathcal{C}^{\perp}$. This decomposition allows decoupling the effect of the generalized consensus steps (38) as follows :
[…]
where $MX_{\mathbin{\!/\mkern-5.0mu/\!}}=X_{\mathbin{\!/\mkern-5.0mu/\!}}$ because $X_{\mathbin{\!/\mkern-5.0mu/\!}}\in\mathcal{C}$ and $M\in\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$.
Theorem 4 ($\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ interpolation constraints).
Let $I$ be a set of $t$ indices and $\lambda^{-},\lambda^{+}\in(-1,1)$, $\lambda^{-}\leq\lambda^{+}$. A set of pairs of points $\{(\mathbf{x}^{k},\mathbf{y}^{k})\}_{k\in I}$ is $\mathcal{M}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$-interpolable if, and only if,